PLEASE USE EXCEL.

SCHOOL    FINAL    GRADE
LAVC    85    A
LAVC    46    D
LAVC    50    D
LAVC    69    C
LAVC    50    D
LAVC    28    F
LAVC    72    C
LAVC    46    F
LAVC    75    C
LAVC    73    C
LAVC    82    B
LAVC    35    F
LAVC    78    A
LAVC    73    C
LAVC    73    C
LAVC    73    C
LAVC    87    B
LAVC    45    F
LAVC    32    F
LAVC    75    A
LAVC    47    C
LAVC    69    C
LAVC    52    C
LAVC    80    A
LAVC    68    C
LAVC    75    C
LAVC    64    B
LAVC    64    B
LAVC    55    F
LAVC    62    D
LAVC    56    F
LAVC    78    B
LAVC    43    F
LAVC    72    C
LAVC    64    C
LAVC    71    C
LAVC    53    F
LAVC    45    A
LAVC    56    C
LAVC    59    A
LAVC    81    A
LAVC    60    C
LAVC    78    B
LAVC    68    A
LAVC    42    F
LAVC    54    A
LAVC    38    D
LAVC    40    C
LAVC    39    F
LAVC    27    F
LAVC    59    A
LAVC    64    A
LAVC    44    B
LAVC    53    A
LAVC    46    C
LAVC    34    F
LAVC    58    A
LAVC    33    C
LAVC    58    D
MONROE HS    32    C
MONROE HS    38    C
MONROE HS    37    C
MONROE HS    50    B
MONROE HS    69    B
MONROE HS        F
MONROE HS    50    A
MONROE HS    38    C
MONROE HS    57    A
MONROE HS    43    C
MONROE HS    65    A
MONROE HS          F
MONROE HS    37    D
MONROE HS    30    C
MONROE HS    62    A
MONROE HS    57    B
MONROE HS    38    C
MONROE HS    34    F
MONROE HS    30    F
MONROE HS    32    D
MONROE HS    30    C
MONROE HS    79    C
MONROE HS    44    B
MONROE HS    72    C
MONROE HS    57    A
MONROE HS    47    C
MONROE HS    49    C
MONROE HS    35    D
MONROE HS    31    F
MONROE HS    55    A
MONROE HS    31    C
MONROE HS    49    A
MONROE HS    45    C
MONROE HS    45    D
MONROE HS    45    D
MONROE HS    39    C
MONROE HS    31    F
ARLETA HS    39    C
ARLETA HS    32    C
ARLETA HS    33    F
ARLETA HS    39    F
ARLETA HS          F
ARLETA HS    49    A
ARLETA HS    58    A
ARLETA HS    66    D
ARLETA HS          F
ARLETA HS    43    D
ARLETA HS    71    C
ARLETA HS          F
ARLETA HS    41    C
ARLETA HS          F
ARLETA HS    38    C
ARLETA HS    40    C
ARLETA HS    31    D
ARLETA HS    37    C
ARLETA HS    89    A
ARLETA HS    48    B
ARLETA HS    62    A
ARLETA HS    40    C
MONROE HS    53    B
MONROE HS    53    B
MONROE HS    39    D
MONROE HS    61    A
MONROE HS    53    F
MONROE HS    47    C
MONROE HS    65    A
MONROE HS    23    F
MONROE HS    36    F
MONROE HS    56    F
MONROE HS    45    C
MONROE HS    40    C
MONROE HS    44    C
MONROE HS    37    D
MONROE HS    20    F
MONROE HS    32    F
MONROE HS    34    C
LAVC    18    F
LAVC    48    B
LAVC    34    D
LAVC    71    C
LAVC    42    B
LAVC    47    A
LAVC    48    B
LAVC    25    D
LAVC    59    A
LAVC    37    C
LAVC    44    B
LAVC    53    A
LAVC    73    C
LAVC    84    B
LAVC    17    F
LAVC    56    D
LAVC    34    D
LAVC    77    C
LAVC    41    C
LAVC    39    B
LAVC    58    A
ARLETA HS    59    A
ARLETA HS    49    A
ARLETA HS    43    D
ARLETA HS    46    B
ARLETA HS    53    B
ARLETA HS    91    B
ARLETA HS    53    B
ARLETA HS    39    D
ARLETA HS    70    A
ARLETA HS    83    B
ARLETA HS    37    F
ARLETA HS    44    C
ARLETA HS    51    C
ARLETA HS    85    B
MONROE HS    60    A
MONROE HS    51    B
MONROE HS    50    C
MONROE HS    60    A
MONROE HS    72    C
MONROE HS    44    F
MONROE HS    69    C
MONROE HS    78    C
MONROE HS    34    F
MONROE HS    73    C
MONROE HS    56    A
LAVC    75    C
LAVC    71    C
LAVC    37    D
LAVC    51    C
LAVC    80    C
LAVC    78    C
LAVC    45    D
LAVC    72    C
LAVC    36    F
LAVC    73    C
LAVC    49    C
LAVC    55    A
LAVC    53    A
LAVC    28    F
LAVC    37    C
LAVC    63    A
LAVC    41    C
LAVC    51    C
LAVC    51    D
LAVC    44    C
LAVC    55    A
LAVC    39    F
LAVC    58    D
LAVC    46    C
LAVC    55    A
LAVC    60    B
LAVC    77    C
LAVC    67    A
LAVC    31    D

What is the scale of measurement for the FINAL variable? Provide the appropriate statistics of central tendency and variability for the FINAL variable and the graph which is most appropriate for the display of those statistics. Summarize in one paragraph what the outut shows.
Answer #1

FINAL variable is discrete..it doe not contain integer.

A discrete variable is a variable with finite range.they are usually integer counts.

The scale of measurement is interval

descriptive statistics are :

mean=sum of values/total values=10666/204=52.28

std deviation=16.19

Graph which is appropriate is box plot.

52.28431373the fivenumber summary is

minimum value=17

Q1=39

Q2=50

Q3=64

maximum value=91

there are no outliers seen from boxplot.

IQR=inter quartile range=Q3-Q1=64-39=25

mean value is higher than median.

Distribution is positively skewed.